OpenStudy (anonymous):

the solution set sqrt3x+16=x+2 is (only 3x+16 is squared)

OpenStudy (anonymous):

x=1/8(67+ sqroot393

OpenStudy (anonymous):

is the left side of the equation square root of (3x+16) or is it (3x+16)^2 ?

OpenStudy (anonymous):

left side is sqrt(3x+16)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if thats the case then you'll have to solve the radical equation by squaring both sides to get 3x+16=x^2+4x+4. getting this in standard form gives x^2+x-15=0. this cannot be solved by factoring so you'll have to use the quadratic formula or complete the square...